Use the following price and quantity information is this table to answer the following questions. Just list the mathematical results to each question in the space below in SNoodle and show your work on your scratch paper 2005 Pre Price 0.00 40 Calculate the Nominal GDP for 2010 and 2011.
Answer #1

Nominal GDP in 2010 = (200)(45) +(2.25)(2200) +(52)(310)

= 9000 + 4950+ 16120

= $ 30070

Nominal GDP in 2011 = (150)(50) + (2.40)(2300) + (55)(350)

= 7500 + 5520+ 19250

= $ 32270